### v1

# Seyed Ali Farokh

Abstract
This paper presents our submissions to the Iranian division of the Text-dependent Speaker Verification Challenge (TdSV) 2024. TdSV aims to determine if a specific phrase was spoken by a target speaker. We developed two independent subsystems based on pre-trained models: For phrase verification, a phrase classifier rejected incorrect phrases, while for speaker verification, a pre-trained ResNet293 with domain adaptation extracted speaker embeddings for computing cosine similarity scores. In addition, we evaluated Whisper-PMFA, a pre-trained ASR model adapted for speaker verification, and found that, although it outperforms randomly initialized ResNets, it falls short of the performance of pre-trained ResNets, highlighting the importance of large-scale pre-training. The results also demonstrate that achieving competitive performance on TdSV without joint modeling of speaker and text is possible. Our best system achieved a MinDCF of 0.0358 on the evaluation subset and won the challenge.
Index Terms: Speaker Recognition, Transfer Learning
1 Introduction
Speaker verification (SV) is the process of confirming a person’s identity through their voice. This verification uses one or more enrollment utterances along with a test utterance and can be either text-dependent (TdSV) or text-independent (TiSV). In a TiSV system, the phonetic content of the utterance is unrestricted, allowing for any spoken content. In contrast, TdSV systems verify both the speaker’s identity and the specific phrase spoken. These systems can be either phrase-dependent, where a fixed set of phrases is pre-defined by the system, or phrase-independent, allowing users to customize their own phrases [1]. With the rising demand for voice-based authentication, TdSV has recently gained more attention, as the phonetic content of utterances can be used as passphrases [2], adding an extra layer of security to voice-based access control systems.
Recent advancements in TiSV have largely focused on improving cascaded systems, which typically consist of a front-end for extracting fixed-size speaker-specific features and a backend for computing similarity scores between speaker embeddings [2]. For the backend, cosine similarity and probabilistic linear discriminant analysis (PLDA) [3] are widely used. The front-end is usually a deep speaker classification network that comprises a frame-level subnetwork for initial feature extraction, a temporal pooling layer for computing fixed-size speaker embeddings from variable-length feature sequences, and a final subnetwork for classification. ResNets [4] and ECAPA-TDNNs [5] are among the most widely used architectures for front-ends, and the most common aggregation techniques include temporal statistics pooling [6] and variants of attentive pooling [7, 8]. Also, advanced softmax variants such as additive margin softmax (AM-Softmax) [9] loss and additive angular margin softmax (AAM-Softmax) [10] loss have been proposed to achieve more discriminative speaker embedding spaces. In addition, pre-trained automatic speech recognition (ASR) models have recently been adapted for speaker verification, showing promising results [11, 12, 13].
TiSV architectures can be effectively adapted for TdSV applications by joint modeling of linguistic and speaker information within a single network. One approach is to use multi-task learning to disentangle speaker and text information, often using two distinct subnetworks for speaker classification and phoneme (or phrase) classification [14, 15]. Further, Liu et al. proposed a phoneme-aware attentive pooling [16] method that incorporates frame-level phoneme posteriors into attentive pooling, enhancing the model’s ability to leverage phonetic information.
This paper presents our system submitted to Task 1 of The Iranian division of The Text-dependent Speaker Verification (TdSV) Challenge 2024111Challenge website: https://tdsvc.github.io [17]. In this challenge, model enrollment is done using three enrollment utterances, and each trial consists of a test utterance and a model identifier. Trials fall into one of the following categories:
  • Target Correct (TC): The test utterance belongs to the claimed model, and the speaker utters the correct phrase.
  • Target Wrong (TW): The test utterance belongs to the claimed model, but the speaker utters an incorrect phrase.
  • Imposter Correct (IC): The speaker’s identity does not match the claimed model, but the speaker utters the correct phrase.
  • Imposter Wrong (IW): The speaker’s identity does not match the claimed model, and the speaker utters an incorrect phrase.
Proposed systems should independently process each trial and generate a score that integrates both speaker and phrase verification scores. The generated scores should only accept TC trials. Also, note that IW trials were excluded from the challenge dataset because they can be easily rejected by a phrase verification system and do not provide valuable insights about a system’s overall performance.
Task 1 of TdSV 2024 is phrase-dependent and uses a fixed set of ten phrases—five in Persian and five in English—for creating enrollment and test recordings. To increase the challenge, some test utterances in TW trials were drawn from free-text recordings.
The primary evaluation metric adopted by TdSV 2024 is the normalized minimum Detection Cost Function (MinDCF), as defined in NIST SRE 2008 as a weighted sum of miss and false error probabilities, with $P_{target}=0.01$, $C_{FalseAlarm}=1$, and $C_{Miss}=10$. Additionally, the Equal Error Rate (EER) metric will be reported.
Unlike most previous studies that adopted joint modeling of speaker and text, our proposed system comprises two independent subsystems for phrase and speaker verification, allowing for the use of various pre-trained models for each subsystem. For phrase verification, we fine-tuned a pre-trained cross-lingual speech representation model for bilingual speech recognition in Persian and English, followed by a further fine-tuning stage for phrase classification. This classifier was used to simply reject TW trials. Similarly, we developed several speaker embedding extractors based on pre-trained ResNets and Whisper [18] for our speaker verification system. After rejecting incorrect phrases using our phrase classifier, we obtained final verification scores by computing cosine similarity between test and enrollment embeddings. Moreover, we used score normalization to enhance the performance further. The experimental results indicated the effectiveness of the proposed method, and our top-performing model ranked first on the challenge leaderboard.
The rest of the paper is organized as follows: Section 2 introduces the datasets used in this work. Sections 3 and 4 describe the architecture of our phrase and speaker verification systems, respectively. The experimental results and discussion are given in Section 5, and we conclude in Section 6.
2 Datasets
The DeepMine dataset [19, 20] is the primary source of the training and evaluation data for TdSV 2024. It was collected through crowd-sourcing, and while all participants were native Persian speakers, most contributed to the English portion of the dataset as well. The official TdSV 2024 data for Task 1 includes three subsets: training, development, and evaluation. The training subset consists of 183,431 utterances from 1,620 speakers. Among the utterances, 31,738 are free-text, while the rest were drawn from a fixed set of ten phrases comprising five Persian and five English phrases. The development and evaluation subsets are intended solely for system evaluation and contain 117,348 and 6,464,241 trials, respectively. During evaluation, model enrollment is conducted using three recordings of a specific phrase, and each trial includes a test utterance and a model identifier. The development set is provided to participants for evaluation and parameter tuning before submitting results to the official leaderboard, whereas the evaluation subset is used for the official evaluation of the challenge. In addition to the DeepMine dataset, participants are also allowed to use the following datasets:
  • VoxCeleb 1&2 [21, 22] are two large-scale datasets collected from YouTube videos. They contain over one million recordings from 7,205 celebrities. In this work, only VoxCeleb 1 was used, which includes over 100,000 utterances from 1,251 speakers.
  • LibriSpeech [23] is a standard ASR corpus in US English that comprises approximately 1,000 hours of speech from 2,338 speakers. We only used the train-clean-100 subset of this dataset to train our phrase verification system, which contains about 100 hours of speech.
  • Common Voice [24] is a multilingual speech dataset created from contributions by volunteers worldwide. For this challenge, teams are restricted to using the Persian (Farsi) subset, which contains approximately 363 hours of validated speech from 4,148 speakers [^2] Common Voice 18.0, released on 6/19/2024. To prepare this subset for training our speaker verification systems, we excluded speakers with fewer than 30 recordings. From the remaining speakers with more than 650 recordings, we randomly selected 650 utterances per speaker, resulting in a final dataset with 125,017 utterances from 813 speakers.
The challenge rules prohibit the use of any other public or private data for training.
2.1 Data augmentation
We did not use any augmentation methods in our phrase verification system. However, following the previous successful studies on speaker verification [25], we adopted SoX-based  [^3] https://sourceforge.net/projects/sox speed perturbation by factors of 0.9 and 1.1 to triple the number of speakers during training, followed by an on-the-fly implementation of the following augmentations with a probability of 0.6:
  • Noise addition with MUSAN [26] dataset
  • RIR reverberation with RIRs [27] dataset
  • Gain augment
3 Phrase verification system
Our proposed system for TdSV 2024 consists of two independent subsystems for phrase and speaker verification. The phrase verification system is a classifier that rejects TW trials, while the speaker verification system is responsible for producing similarity scores. Although this system design does not benefit from joint modeling of speaker and text, it greatly simplifies the system development process and allows for the use of various pre-trained models for each subsystem with minimal modifications.
The phrase classifier is an 11-class model trained with standard softmax. The first ten classes correspond to the set of phrases in the challenge, and the final class represents free text (or ”none of the above”). This classifier is built on XLSR [28], a pre-trained cross-lingual speech representation model trained by solving a self-supervised contrastive task, proven to be effective in low-resource languages compared to traditional feature extraction methods. This model takes a raw waveform as input and produces a sequence of features.
Moreover, to improve the model’s ability to extract linguistic features from Persian and English inputs, we initially fine-tuned the XLSR for bilingual speech recognition in Persian and English. During this phase, 30% of the training subsets of Common Voice Farsi and LibriSpeech (train-clean-100) were used, and the model was trained using CTC loss [29] for 40 epochs, with an initial learning rate of 0.001 and an effective batch size of 32.
Finally, for training the classifier, an attention-based pooling layer was added to the fine-tuned XLSR to compute utterance-level feature vectors from frame-level representations $h_{t}$ ($t=1,...,T$):
$e_{t}=W_{1}h_{t}+b_{1},$ (1)
$\alpha_{t}=\frac{\exp(e_{t})}{\sum_{\tau}^{T}\exp(e_{\tau})},$ (2)
$\tilde{h}=\sum_{t}^{T}\alpha_{t}(W_{2}h_{t}+b_{2}),$ (3)
where $e_{t}$ and $\alpha_{t}$ are the attention score and weight, respectively. $\tilde{h}$ refers to the utterance-level feature vector, which is finally fed to a fully connected layer with ReLU activation, followed by a linear classifier. The network was trained using Cross-Entropy loss function for one epoch on the entire training samples of the challenge dataset, with a learning rate of 0.0005 and an effective batch size of 64.
Table 1: Hyper-parameters used across different training stages. BS and LR denote batch size and learning rate.
System Full Training Domain Adaptation
Epoch BS LR Epoch BS LR
S2 - - - 15 32 3e-4
S3 100 64 1e-3 15 32 3e-4
S4 15 64 1e-3 7 28 5e-5
S5 15 64 1e-3 7 28 5e-5
4 Speaker verification system
Our primary goal in Task 1 of TdSV 2024 was to leverage the power of pre-trained models to overcome the challenges of in-domain data scarcity and mitigate the computational cost of training randomly initialized models. We explored two directions for developing speaker verification models and compared their results. In the first approach, we trained several ResNet-based models, which have been widely used as a standard architecture in speaker verification. In the second approach, we studied the performance of pre-trained ASR models adapted for the speaker verification task. More specifically, we employed the Whisper-PMFA [30] method, which involves fine-tuning a pre-trained Whisper model for speaker recognition.
4.1 Training protocol
We trained our models in two stages:
  • Full training (T1): In this stage, models were trained on a combination of out-of-domain data (Common Voice Farsi and VoxCeleb 1) and in-domain (DeepMine) data, totaling 3,684 speakers, to learn robust and generalizable speaker embeddings across different datasets. pre-trained ResNets did not undergo this stage. During this phase, 300 frames of each sample input were randomly selected in each epoch to prevent overfitting and accelerate the training. Moreover, all augmentation methods explained in Section 2.1 were applied. We employed the widely used AAM-Softmax [10] loss with the subcenter method and the Inter-TopK penalty [31] to train our models, with a constant margin and scale of 0.2 and 32, respectively.
  • Domain adaptation (T2): We fine-tuned our models with in-domain data after full training to bridge the domain gap and enhance the performance. During this stage, augmentation methods and the Inter-TopK penalty were removed to avoid domain mismatch, and the number of randomly selected frames was increased from 300 to 600. Also, fine-tuning was done with smaller learning rate values to maintain the generalization abilities of the models.
All models were optimized using SGD with a momentum of 0.9 and a weight decay of 1e-4. We also utilized an exponential decay scheduler with a minimum learning rate of 5e-5 for T1 and 1e-6 for T2. Other training hyper-parameters are listed in Table 1. Note that gradient accumulation was used to achieve the target effective batch size when GPU memory was limited. The dimensionality of speaker embeddings was set to 256 across all models. All experiments were conducted on a single Nvidia RTX 3090 GPU using the WeSpeaker toolkit [32, 33].
4.2 ResNet
ResNet [4] is a widely used architecture for speaker recognition that has performed excellently in previous speaker verification challenges [25]. Consequently, many open-source implementations and pre-trained models have been publicly released based on this architecture. Trained on large-scale datasets like VoxCeleb 1&2, these pre-trained models can provide a robust starting point for training speaker recognition models on other datasets by improving their generalization and speeding up the convergence.
During the challenge period, we submitted three systems based on bottleneck-block-based ResNet, all of which adopted the temporal statistics pooling [6] mechanism to aggregate variable-length sequence features to utterance-level embeddings. The first system (S1) was a pre-trained ResNet34 without domain adaptation, while the second one (S2) was a pre-trained ResNet293 that underwent domain adaptation. Finally, we applied both training stages to a randomly initialized ResNet152 to obtain our last ResNet-based system (S3).
Table 2: Phrase verification results on TC-vs-TW trials
Subset MinDCF0.01 EER(%)
development 0.0000 0.00
evaluation 0.0003 0.01
Table 3: Results on the development and evaluation sets
System Architecture Training Stages Development Evaluation
MinDCF0.01 EER(%) MinDCF0.01 EER(%)
S1 ResNet34 0.0614 1.3938 0.0784 1.7390
S2 ResNet293 T2 0.0225 0.8733 0.0376 1.1080
S3 ResNet152 T1 + T2 0.0191 0.6757 0.0764 2.3444
S4 Whisper-PMFA T1 + T2 0.0163 0.6121 0.0584 2.0410
S5 Whisper-PMFA T1 + T2 0.0161 0.6126 0.0583 2.0445
[…]
4.3 Whisper-PMFA
Building on the successful use of pre-trained ASR models in speaker verification [11, 12], Whisper-PMFA (Partial Multi-Scale Feature Aggregation using Whisper) [30] was recently introduced to leverage the capabilities of Whisper, a large-scale multilingual ASR model based on transformers [34]. Whisper-PMFA adapts Whisper for speaker verification by selectively concatenating frame-level outputs from specific transformer layers, rather than aggregating features from all layers. This approach not only reduces computational overhead but also enhances performance by minimizing the integration of irrelevant information from lower-impact layers.
Inspired by this, we studied the performance of Whisper-PMFA in this challenge. Since Whisper was not trained for the speaker recognition task, we applied both training stages to Whisper-PMFA. Additionally, before the full training stage, we froze the Whisper parameters and fine-tuned the model for five epochs to prevent updating the pre-trained model in the wrong direction due to the random initialization of newly added components.
We submitted two Whisper-PMFA-based systems to this challenge (S5 and S6), with the only difference being the AAM-Softmax margin during the domain adaptation phase, which was set to 0.35 and 0.2 for S5 and S6, respectively.
4.4 Feature extraction
80-dimensional log Mel filter bank energies with a 25ms window and 10ms frame shift were extracted for our ResNet-based models. Voice activity detection (VAD) was not applied, and all features were mean-normalized. Likewise, 80-dimensional log magnitude Mel spectrograms consistent with the pre-trained Whisper were utilized for training Whisper-PMFA.
4.5 Backend
Speaker embeddings were extracted from the final fully connected layer of the models, and cosine similarity was used to compute scores. Since model enrollment is done using three utterances in this challenge, we used the average of embedding vectors of each model during scoring.
Afterward, AS-Norm [35] was used for score normalization, using 1,620 cohorts obtained from speaker-wise averaging of all embeddings in the training subset of the challenge dataset. Top-300 most similar scores were selected to compute the mean and standard deviation for normalization.
Finally, we adopted score fusion by averaging single-system scores to further improve performance.
5 Results
Table 3 shows the evaluation results of our single and fusion systems on the development and evaluation subsets of the challenge after applying AS-Norm and rejecting TW trials. The results indicate that the Whisper-PMFA method outperforms the widely used ResNet architecture, conforming to the findings of previous studies on the effectiveness of adapting pre-trained ASR models for speaker verification. However, it can be observed from the results that the ResNets pre-trained on approximately twice the data (VoxCeleb 1&2) can surpass Whisper-PMFA after a simple domain adaptation stage, which highlights the importance of large-scale pre-training in improving the generalization ability of speaker verification models.
We also report the MinDCF and EER of the proposed phrase verification system on TC-vs-TW trials of the development and evaluation subsets (Table 2). According to the results, our phrase verification system demonstrates a near-optimal performance on this task.
5.1 Ablation study
We conducted an ablation study on our Whisper-PMFA system (S5). The development set of the challenge dataset was used as our evaluation benchmark. We can observe from the results (Table 4) that the domain adaptation phase improved the MinDCF from 0.0234 to 0.0177. Also, a further improvement of MinDCF to 0.0161 was achieved after applying AS-Norm.
Table 4: Ablation study
Methods Development
MinDCF0.01 EER(%)
Whisper-PMFA (T1) 0.0234 0.9253
+ Domain adaptation (T2) 0.0177 0.6273
++ AS-Norm 0.0161 0.6126
6 Conclusion
In this report, we described our system for Task 1 of the Iranian division of the TdSV challenge 2024. Our phrase verification system was a classifier built on top of XLSR, and our speaker verification models were based on Whisper-PMFA and pre-trained ResNets. The experimental results show that, in this task, the performance of Whisper-PMFA is inferior to a pre-trained ResNet293 with domain adaptation, even though it can outperform the ResNet architecture with random initialization and the same training data. The results also demonstrate the possibility of developing competitive text-dependent speaker verification systems using independent pre-trained components. Finally, our best system was a fusion of multiple ResNet- and Whisper-PMFA-based models, which achieved a MinDCF of 0.0358 on the evaluation subset and won the challenge.